omework G Heip Sav a. The company completed consulting work b. The company completed commission work for a client and sent a bill for $5 600 to be received within 30 dsys c. The company paid an assistant $2,200 cash as wages for the period d. The company collected $2,800 cash as a partial payment for the amount owed by the client in transaction e. The company paid $1,020 cash for this periods cleaning services for a client and immediately collected $7100 cash earned Required: Enter the impact of each transaction on individual items of the accounting equation (Enter decreases to account bolances with minus sign) Equity Revenue Liabilities Assets CommonD Accounts Accounts Payable - Dividends Cash Receivable b. Bal. Bal. d. Bal. 扫扫: Next > Bal. くPrgy 50f 10 search
Answer #1

Solution :-

Cash + Accounts Receivable = Accounts Payable + Common Stock - Dividends + Revenue - Expense
a 7100 + = + - + 7100 -
b + 5600 = + - + 5600 -
Bal 7100 + 5600 = + - + 12700 -
c -2200 + = + - + - 2200
Bal 4900 + 5600 = + - + 12700 - 2200
d 2800 + -2800 = + - + -
Bal 7700 + 2800 = + - + 12700 - 2200
e -1020 + = + - + - 1020
Bal 6680 + 2800 = + - + 12700 - 3220